OpenStudy (anonymous):

Find the value of x that makes the statement true. 3/5 x-5+x-4=-1

OpenStudy (solomonzelman):

\(\LARGE\color{midnightblue}{ \rm \frac{3}{5}x-5+x-4=-1 }\) \(\LARGE\color{midnightblue}{ \rm \frac{3}{5}x+x-9=-1 }\) \(\LARGE\color{midnightblue}{ \rm \frac{3}{5}x+x=8 }\) \(\LARGE\color{midnightblue}{ \rm \frac{3x}{5}+x=8 }\) \(\LARGE\color{midnightblue}{ \rm \frac{3x}{5}+\frac{5x}{5}=\frac{40}{5} }\) \(\LARGE\color{midnightblue}{ \rm 8x=40}\)